OpenStudy (anonymous):

A group of 10 girls and 6 boys from the sophomore class board are in charge of the 2005 prom. In how many ways can they elect a President, Vice-President and Treasurer?

OpenStudy (anonymous):

First choose 3 people from the 16.That is done in 16C3 ways. Now out of the three chosen there are 3! ways in which they can be made President,Vice President and treasurer. So the total number of ways is 16C3*3!= 16*15*14*6/6=1960

OpenStudy (anonymous):

or just compute \[16\times 15\times 14\]

OpenStudy (anonymous):

16 choices for president then 15 for vice president then 14 for treasurer

OpenStudy (anonymous):

which is 3360, not the same as the answer above

OpenStudy (anonymous):

@beekzebub your answer is correct, but you number is wrong i think. also there is no reason to do all that work, especially multiplying by 6 and dividing by 6.

OpenStudy (anonymous):

yea, sorry 3360